# On-Policy Policy Gradient Reinforcement Learning Without On-Policy Sampling

### v2

@@ Abstract
On-policy reinforcement learning (rl) algorithms perform policy updates using i.i.d. trajectories collected by the current policy. However, after observing only a finite number of trajectories, on-policy sampling may produce data that fails to match the expected on-policy data distribution. This sampling error leads to noisy updates and data inefficient on-policy learning. Recent work in the policy evaluation setting has shown that non-i.i.d., off-policy sampling can produce data with lower sampling error than on-policy sampling can produce [Zhong et al., 2022]. Motivated by this observation, we introduce an adaptive, off-policy sampling method to improve the data efficiency of on-policy policy gradient algorithms. Our method, Proximal Robust On-Policy Sampling (props), reduces sampling error by collecting data with a behavior policy that increases the probability of sampling actions that are under-sampled with respect to the current policy. We empirically evaluate props on both continuous-action MuJoCo benchmark tasks as well discrete-action tasks and demonstrate that (1) props decreases sampling error throughout training and (2) improves the data efficiency of on-policy policy gradient algorithms.
@@ 1 Introduction
One of the most widely used classes of reinforcement learning (rl) algorithms is the class of on-policy policy gradient algorithms. These algorithms use gradient ascent on the parameters of a parameterized policy so as to increase the probability of observed actions with high expected returns under the current policy. The gradient is commonly estimated using the Monte Carlo estimator, an average computed over i.i.d. samples of trajectories from the current policy. The Monte Carlo estimator is consistent and unbiased; as the number of sampled trajectories increases, the empirical distribution of trajectories converges to the true distribution under the current policy, and thus the empirical gradient converges to the true gradient. However, the expense of environment interactions forces us to work with finite samples. Thus, the empirical distribution of the trajectories often differs from the desired on-policy data distribution. This sampling error produces inaccurate gradient estimates, resulting in noisy policy updates, slower learning, and potentially convergence to suboptimal policies. With i.i.d. on-policy sampling, the only way to reduce sampling error is to collect more data.
@@
Since on-policy sampling is so widely used to produce on-policy data, on-policy sampling is often taken to be an essential feature of data collection for on-policy learning [Silver, 2015, Achiam, 2018]. However, on-policy sampling is not explicitly required for consistent policy updates; on-policy learning requires on-policy data – data whose state-conditioned empirical distribution of actions matches that of the current policy. On-policy sampling is a straightforward way to acquire on-policy data, though we can obtain such data more efficiently without on-policy sampling. To better illustrate this concept, consider an MDP with two discrete actions A and B, and suppose the current policy $\pi$ places equal probability on both actions in some state $s$. When following $\pi$, after 10 visits to $s$, we may observe A 2 times and B 8 times rather than the expected 5 times. Alternatively, if we adaptively select the most under-sampled action upon every visit to $s$, we will observe each action an equal number of times. The first scenario illustrates on-policy sampling but not on-policy data; the second scenario uses off-policy sampling yet produces on-policy data.
These observations raise the following question: can on-policy policy gradient algorithms learn more efficiently using on-policy data acquired without on-policy sampling? Recently, Zhong et al. [2022] showed that adaptive, off-policy sampling can yield data that more closely matches the on-policy distribution than data produced by i.i.d. on-policy sampling. However, this work was limited to the policy evaluation setting in which the on-policy distribution remains fixed. Turning from evaluation to control poses the challenge of a continually changing current policy.
@@
In this work, we address this challenge and show for the first time that on-policy policy gradient algorithms are more data-efficient learners when they use on-policy data acquired with adaptive, off-policy sampling. We introduce a method, Proximal Robust On-Policy Sampling (props) , that adaptively corrects sampling error in previously collected data by increasing the probability of sampling actions that are under-sampled with respect to the current policy. Figure 1 provides an overview of props. We empirically evaluate props on continuous-action MuJoCo benchmark tasks as well as discrete action tasks and show that (1) props reduces sampling error throughout training and (2) that doing so improves the data efficiency of on-policy policy gradient algorithms. In summary, our contributions are
@@
    We introduce a scalable adaptive sampling algorithm that reduces sampling error in on-policy data collection.
@@
    We demonstrate empirically that our method improves the data efficiency of on-policy policy gradient algorithms and increases the fraction of training runs that converge to high-return polices.
@@
    Building off of the theoretical foundation laid by Zhong et al. [2022], this work improves the rl community’s understanding of a nuance in the on-policy vs off-policy dichotomy: on-policy learning requires on-policy data, not on-policy sampling.
Figure 1: An overview of props for on-policy policy gradient learning. Rather than collecting data ${\mathcal{D}}$ via on-policy sampling from the agent’s current policy $\pi_{\bm{\theta}}$, we collect data with a separate data collection policy $\pi_{\bm{\phi}}$ that we continually adapt to reduce sampling error in ${\mathcal{D}}$ with respect to the agent’s current policy.
@@ 2 Related Work
Our work focuses on data collection in rl. In rl, data collection is often framed as an exploration problem, focusing on how an agent should explore its environment to efficiently learn an optimal policy. Prior rl works have proposed several exploration-promoting methods such as intrinsic motivation [Pathak et al., 2017, Sukhbaatar et al., 2018], count-based exploration [Tang et al., 2017, Ostrovski et al., 2017], and Thompson sampling [Osband et al., 2013, Sutton and Barto, 2018]. In contrast, our objective is to learn from the on-policy data distribution; we use adaptive data collection to more efficiently obtain this data distribution.
@@
Prior works have used adaptive off-policy sampling to reduce sampling error in the policy evaluation subfield of rl. Most closely related is the work of Zhong et al. [2022] who first proposed that adaptive off-policy sampling could produce data that more closely matches the on-policy distribution than on-policy sampling could produce. Mukherjee et al. [2022] use a deterministic sampling rule to take actions in a particular proportion. Other bandit works use a non-adapative exploration policy to collect additional data conditioned on previously collected data [Tucker and Joachims, 2022, Wan et al., 2022, Konyushova et al., 2021]. Since these works only focus on policy evaluation, they do not have to contend with a changing on-policy distribution as our work does for the control setting.
@@
Several prior works propose importance sampling methods [Precup, 2000] to reduce sampling error without further data collection. In the rl setting, Hanna et al. [2021] showed that reweighting off-policy data according to an estimated behavior policy can correct sampling error and improve policy evaluation. Similar methods have been studied for temporal difference learning [Pavse et al., 2020] and policy evaluation in the bandit setting [Li et al., 2015, Narita et al., 2019]. Conservative Data Sharing [Yu et al., 2021] reduces sampling error by selectively integrating offline data from multiple tasks. Our work instead focuses on using additional data collection to reduce sampling error.
@@
As we will discuss in Section 5, the method we introduce permits data collected in one iteration of policy optimization to be re-used in future iterations rather than discarded as typically done by on-policy algorithms. Prior work has attempted to avoid discarding data by combining off-policy and on-policy updates with separate loss functions or by using alternative gradient estimates [Wang et al., 2016, Gu et al., 2016, 2017, Fakoor et al., 2020, O’Donoghue et al., 2016, Queeney et al., 2021]. In contrast, our method modifies the sampling distribution at each iteration so that the entire data set of past and newly collected data matches the expected distribution under the current policy.
@@ 3.1 Reinforcement Learning
We formalize the rl environment as a finite horizon Markov decision process (MDP) [Puterman, 2014] $({\mathcal{S}},{\mathcal{A}},p,r,d_{0},\gamma)$ with state space ${\mathcal{S}}$, action space ${\mathcal{A}}$, transition dynamics $p:{\mathcal{S}}\times{\mathcal{A}}\times{\mathcal{S}}\to[0,1]$, reward function $r:{\mathcal{S}}\times{\mathcal{A}}\to\mathbb{R}$, initial state distribution $d_{0}$, and reward discount factor $\gamma\in[0,1)$. The state and action spaces may be discrete or continuous. We write $p(\cdot\mid{\bm{s}},{\bm{a}})$ to denote the distribution of next states after taking action ${\bm{a}}$ in state ${\bm{s}}$. We consider stochastic policies $\pi_{\bm{\theta}}:{\mathcal{S}}\times{\mathcal{A}}\to[0,1]$ parameterized by ${\bm{\theta}}$, and we write $\pi_{\bm{\theta}}({\bm{a}}|{\bm{s}})$ to denote the probability of sampling action ${\bm{a}}$ in state ${\bm{s}}$ and $\pi_{\bm{\theta}}(\cdot|{\bm{s}})$ to denote the probability distribution over actions in state ${\bm{s}}$. We additionally let $d_{\pi_{\bm{\theta}}}:{\mathcal{S}}\times{\mathcal{A}}\to[0,1]$ denote the state-action visitation distribution, the distribution over state-action pairs induced by following $\pi_{\bm{\theta}}$. The rl objective is to find a policy that maximizes the expected sum of discounted rewards, defined as:
$\begin{split}J({\bm{\theta}})=\mathbb{E}_{({\bm{s}},{\bm{a}})\sim d_{\pi_{\bm{\theta}}}}\left[\sum\nolimits_{t=0}^{H}\gamma^{t}r({\bm{s}}_{t},{\bm{a}}_{t})\right],\end{split}$ (1)
where the horizon $H$ is the random variable representing the time-step when an episode ends. Throughout this paper, we refer to the policy used for data collection as the behavior policy and the policy trained to maximize its expected return as the target policy.
@@ 3.2 On-Policy Policy Gradient Algorithms
Policy gradient algorithms are one of the most widely used methods in rl. These methods perform gradient ascent over policy parameters to maximize an agent’s expected return $J({\bm{\theta}})$ (Eq. 1). The gradient of the $J({\bm{\theta}})$ with respect to ${\bm{\theta}}$, or policy gradient, is often given as:
$\nabla_{\bm{\theta}}J({\bm{\theta}})=\mathbb{E}_{({\bm{s}},{\bm{a}})\sim d_{\pi_{\bm{\theta}}}}\left[A^{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}({\bm{a}}|{\bm{s}})\right],$ (2)
@@
where $A^{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})$ is the advantage of choosing action ${\bm{a}}$ in state ${\bm{s}}$ and following $\pi_{\bm{\theta}}$ thereafter. In practice, the expectation in Eq. 2 is approximated with Monte Carlo samples collected from $\pi_{\bm{\theta}}$ and an estimate of $A^{\pi_{\bm{\theta}}}$ used in place of the true advantages [Schulman et al., 2016]. After updating the policy parameters with this estimated gradient, the previously collected trajectories ${\mathcal{D}}$ become off-policy with respect to the updated policy. To ensure gradient estimation remains unbiased, on-policy algorithms discard historic data after each update and collect new data with the updated policy.
@@
This foundational idea of policy learning via stochastic gradient ascent was first proposed by Williams [Williams, 1992] under the name reinforce. Since then, a large body of research has focused on developing more scalable policy gradient methods [Kakade, 2001, Schulman et al., 2015, Mnih et al., 2016, Espeholt et al., 2018, Lillicrap et al., 2015, Haarnoja et al., 2018]. Arguably, the most successful variant of policy gradient learning is proximal policy optimization (ppo[Schulman et al., 2017], the algorithm of choice in several high-profile success stories [Berner et al., 2019, Akkaya et al., 2019, Vinyals et al., 2019]. Rather than maximizing the standard rl objective (Eq. 1), ppo maximizes a surrogate objective:
@@
$\begin{split}{\mathcal{L}}_{\text{{ppo}{}}}&({\bm{s}},{\bm{a}},{\bm{\theta}},{\bm{\theta}}_{\text{old}})=\min(g({\bm{s}},{\bm{a}},{\bm{\theta}},{\bm{\theta}}_{\text{old}})A^{\pi_{{\bm{\theta}}_{\text{old}}}}({\bm{s}},{\bm{a}}),\\ &\mathtt{clip}(g({\bm{s}},{\bm{a}},{\bm{\theta}},{\bm{\theta}}_{\text{old}}),1-\epsilon,1+\epsilon)A^{\pi_{{\bm{\theta}}_{\text{old}}}}({\bm{s}},{\bm{a}})),\end{split}$ (3)
@@
where ${\bm{\theta}}_{\text{old}}$ denotes the policy parameters prior to the update, $g({\bm{s}},{\bm{a}},{\bm{\theta}},{\bm{\theta}}_{\text{old}})$ is the policy ratio $g({\bm{s}},{\bm{a}},{\bm{\theta}},{\bm{\theta}}_{\text{old}})=\frac{\pi_{\bm{\theta}}({\bm{a}}|{\bm{s}})}{\pi_{{\bm{\theta}}_{\text{old}}}({\bm{a}}|{\bm{s}})}$, and the $\mathtt{clip}$ function with hyperparameter $\epsilon$ clips $g({\bm{s}},{\bm{a}},{\bm{\theta}},{\bm{\theta}}_{\text{old}})$ to the interval $[1-\epsilon,1+\epsilon]$. The first term inside the minimum of ${\mathcal{L}}_{\text{{ppo}{}}}$ is the conservative policy iteration (CPI) objective [Kakade and Langford, 2002]. The second term clips the CPI objective to disincentivizes large changes to $\pi_{\bm{\theta}}({\bm{a}}|{\bm{s}})$. In contrast to other policy gradient algorithms which perform a single gradient update per data sample to avoid destructively large weight updates, ppo’s clipping mechanism allows the agent to perform multiple epochs of minibatch policy updates.

### v3

@@ Abstract
On-policy reinforcement learning (RL) algorithms are typically characterized as algorithms that perform policy updates using i.i.d. trajectories collected by the agent’s current policy. However, after observing only a finite number of trajectories, such on-policy sampling may produce data that fails to match the expected on-policy data distribution. This sampling error leads to high-variance gradient estimates that yield data-inefficient on-policy learning. Recent work in the policy evaluation setting has shown that non-i.i.d., off-policy sampling can produce data with lower sampling error w.r.t. the expected on-policy distribution than on-policy sampling can produce (Zhong et al., 2022). Motivated by this observation, we introduce an adaptive, off-policy sampling method to reduce sampling error during on-policy policy gradient RL training. Our method, Proximal Robust On-Policy Sampling (PROPS), reduces sampling error by collecting data with a behavior policy that increases the probability of sampling actions that are under-sampled w.r.t. the current policy. We empirically evaluate PROPS on continuous-action MuJoCo benchmark tasks as well as discrete-action tasks and demonstrate that (1) PROPS decreases sampling error throughout training and (2) increases the data efficiency of on-policy policy gradient algorithms.
@@ 1 Introduction
One of the most widely used classes of reinforcement learning (RL) algorithms is the class of on-policy policy gradient algorithms. These algorithms use gradient ascent on the parameters of a parameterized policy to increase the probability of observed actions with high expected returns under the current policy. The gradient is commonly estimated using the Monte Carlo estimator, an average computed over i.i.d. trajectories sampled from the current policy. The Monte Carlo estimator is consistent and unbiased; as the number of sampled trajectories increases, the empirical distribution of data converges to the expected distribution under the current policy, and thus the estimated gradient converges to the true gradient. However, the expense of environment interaction forces us to work with finite sample sizes. Thus, the empirical distribution of data often differs from the desired on-policy data distribution, a mismatch we call sampling error. Sampling error causes inaccurate gradient estimates, resulting in high-variance policy updates, slower learning, and potentially convergence to suboptimal policies.
@@
With i.i.d. on-policy sampling, the only way to reduce sampling error is to collect more data. Alternatively, we can reduce sampling error more efficiently using adaptive, off-policy sampling. To illustrate this approach, consider an MDP with two discrete actions A and B, and suppose the current policy $\pi$ places equal probability on both actions in some state $s$. When following $\pi$, after 10 visits to $s$, we will observe both actions 5 times in expectation. Now suppose that after the first 9 visits to $s$, we actually observe A 4 times and B 5 times. If we sample an action from $\pi$ upon our next visit to $s$, we may sample B, and our data will not match the expected on-policy distribution. Alternatively, if we select the under-sampled action A with probability 1, we will observe each action 5 times, making the aggregate data match the on-policy distribution even though this final action was sampled off-policy. The first scenario illustrates on-policy sampling with sampling error; the second scenario uses adaptive, off-policy sampling to reduce sampling error.
Recently, Zhong et al. (2022) introduced an adaptive, off-policy sampling method (ROS) that can produce data that more closely matches the on-policy distribution than data acquired through i.i.d. on-policy sampling. However, this work was limited to low-dimensional policy evaluation tasks. Moreover, ROS required large batches of data to reduce sampling error—approximately 5000 samples on tasks like CartPole-v1 (Brockman et al., 2016)—and struggled to reduce sampling error on tasks with continuous actions. To make adaptive sampling practical for data-efficient RL, we need methods that can reduce sampling error in high-dimensional continuous-action tasks while using the smaller batch sizes typically used in RL. These observations raise the following question: Can reducing sampling error through adaptive sampling increase the data-efficiency of on-policy policy gradient methods?
@@
In this work, we address these challenges and show for the first time that on-policy policy gradient algorithms are more data-efficient learners when they use adaptive, off-policy sampling to reduce sampling error w.r.t. the on-policy distribution. Our method, Proximal Robust On-Policy Sampling (PROPS), adaptively corrects sampling error in previously collected data by increasing the probability of sampling actions that are under-sampled with respect to the current policy. Fig. 1 provides an overview of PROPS. We empirically evaluate PROPS on continuous-action MuJoCo benchmark tasks and show that (1) PROPS reduces sampling error throughout training and (2) increases the data efficiency of on-policy policy gradient algorithms. In summary, our contributions are:
@@
    We introduce PROPS, a scalable adaptive sampling algorithm for on-policy policy gradient learning that reduces sampling error w.r.t. the agent’s current policy.
@@
    We demonstrate empirically that PROPS reduces sampling error more efficiently than on-policy sampling and ROS.
@@
    We demonstrate empirically that reducing sampling error via PROPS increases data efficiency in on-policy policy gradient RL.
Figure 1: An overview of PROPS. Rather than collecting data $\mathcal{D}$} via on-policy sampling from the agent’s current policy $\pi_{\bm{\theta}}$, we collect data with a separate data collection policy $\pi_{\bm{\phi}}$ that we continually adapt to reduce sampling error in $\mathcal{D}$} with respect to the agent’s current policy.
@@ 2 Related Work
Data Collection and Exploration. Our work focuses on data collection in RL. In RL, data collection is often framed as an exploration problem, focusing on how an agent should explore its environment to efficiently learn an optimal policy. Prior works have proposed exploration-promoting methods such as intrinsic motivation (Pathak et al., 2017; Sukhbaatar et al., 2018), count-based exploration (Tang et al., 2017; Ostrovski et al., 2017), and Thompson sampling (Osband et al., 2013; Sutton and Barto, 2018). Rather than adjusting the agent’s sampling distribution to promote exploration, our objective is to learn from the on-policy data distribution; we use adaptive data collection to more efficiently obtain this data distribution.
@@
Adaptive Sampling for Sampling Error Reduction. Prior works have used adaptive off-policy sampling to reduce sampling error in the policy evaluation subfield of RL. Most closely related is the work of Zhong et al. (2022), who first proposed that adaptive off-policy sampling could produce data that more closely matches the on-policy distribution than on-policy sampling could produce. Mukherjee et al. (2022) use a deterministic sampling rule to take actions in a particular proportion. Other bandit works use a non-adaptive exploration policy to collect additional data conditioned on previously collected data (Tucker and Joachims, 2022; Wan et al., 2022; Konyushova et al., 2021). Since these works only focus on policy evaluation, they do not have to contend with a changing on-policy distribution as our work does for the control setting.
@@
Importance Sampling for Sampling Error Reduction. Several prior works propose importance sampling methods (Precup, 2000) to reduce sampling error without further data collection. In the RL setting, Hanna et al. (2021) showed that reweighting off-policy data according to an estimated behavior policy can correct sampling error and improve policy evaluation. Liu et al. (2024) compute a globally optimal behavior policy using an offline dataset. Papini et al. (2024) focused on finding a behavior policy that minimizes the variance of importance-weighted policy gradient estimates. Similar methods have been studied for temporal difference learning (Pavse et al., 2020) and policy evaluation in the bandit setting (Li et al., 2015; Narita et al., 2019). Conservative Data Sharing (Yu et al., 2021) reduces sampling error by selectively integrating offline data from multiple tasks. While these works reduce sampling error by reweighting existing data, our work focuses on whether we can reduce sampling error during data collection. Moreover, works like Liu et al. (2024) require a large amount of offline data (e.g., 1000 trajectories) to accurately compute a behavior policy, but in online on-policy RL, agents typically collect only a few trajectories between policy updates.
@@
On-Policy Learning with Off-Policy Data. As we will discuss in Section 5, the method we introduce permits data collected in one iteration of policy optimization to be re-used in future iterations rather than discarded as typically done by on-policy algorithms. Prior work has attempted to avoid discarding data by combining off-policy and on-policy updates with separate loss functions or by using alternative gradient estimates (Wang et al., 2016; Gu et al., 2016; 2017; Fakoor et al., 2020; O’Donoghue et al., 2016; Queeney et al., 2021). In contrast, our method modifies the sampling distribution at each iteration so that the entire data set of past and newly collected data matches the expected distribution under the current policy.
@@ 3.1 Reinforcement Learning
We formalize the RL environment as an infinite horizon Markov decision process (MDP) (Puterman, 2014) $({\mathcal{S}},{\mathcal{A}},p,r,d_{0},\gamma)$ with state space $\mathcal{S}$}, action space $\mathcal{A}$}, transition dynamics $p:{\mathcal{S}}\times{\mathcal{A}}\times{\mathcal{S}}\to[0,1]$, reward function $r:{\mathcal{S}}\times{\mathcal{A}}\to\mathbb{R}$, initial state distribution $d_{0}$, and reward discount factor $\gamma\in[0,1)$. The state and action spaces may be discrete or continuous. We consider stochastic policies $\pi_{\bm{\theta}}:{\mathcal{S}}\times{\mathcal{A}}\to[0,1]$ parameterized by $\bm{\theta}$}, and we write $\pi_{\bm{\theta}}({\bm{a}}|{\bm{s}})$ to denote the probability of sampling action $\bm{a}$} in state $\bm{s}$} and $\pi_{\bm{\theta}}(\cdot|{\bm{s}})$ to denote the probability distribution over actions in state $\bm{s}$}. We additionally let $d_{\pi_{\bm{\theta}}}:{\mathcal{S}}\times{\mathcal{A}}\to[0,1]$ denote the state-action visitation distribution, the distribution over state-action pairs induced by following $\pi_{\bm{\theta}}$. The RL objective is to find a policy that maximizes the expected sum of discounted rewards, defined as:
$\begin{split}J({\bm{\theta}})=\mathbb{E}_{\tau\sim{\pi_{\bm{\theta}}}}\left[\sum\nolimits_{t=0}^{\infty}\gamma^{t}r({\bm{s}}_{t},{\bm{a}}_{t})\right].\end{split}$ (1)
Throughout this paper, we refer to the policy used for data collection as the behavior policy and the policy trained to maximize its expected return as the target policy.
@@ 3.2 On-Policy Policy Gradient Algorithms
Policy gradient algorithms perform gradient ascent over policy parameters to maximize an agent’s expected return $J({\bm{\theta}})$. The gradient of $J({\bm{\theta}})$ with respect to $\bm{\theta}$}, or policy gradient, is often given as:
$\nabla_{\bm{\theta}}J({\bm{\theta}})=\mathbb{E}_{{\bm{s}}\sim d_{\pi_{\bm{\theta}}},{\bm{a}}\sim{\pi_{\bm{\theta}}}}\left[A^{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}({\bm{a}}|{\bm{s}})\right],$ (2)
@@
where $A^{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})$ is the advantage of choosing action $\bm{a}$} in state $\bm{s}$} and following $\pi_{\bm{\theta}}$ thereafter. [^1] We note that Eq. 2 is a biased estimator of the policy gradient; the expectation is taken over the undiscounted state distribution $d_{\pi_{\bm{\theta}}}$ rather than the discounted state distribution $d^{\gamma}_{\pi_{\bm{\theta}}}$, so Eq. 2 is not the gradient of the discounted objective Eq. 1 (Nota and Thomas, 2019). Nevertheless, since many popular policy gradient algorithms estimate the policy gradient using Eq. 2 (Schulman et al., 2017; Haarnoja et al., 2018; Mnih et al., 2016), we use $d_{\pi_{\bm{\theta}}}$ rather than $d^{\gamma}_{\pi_{\bm{\theta}}}$ in this work. In practice, the expectation in Eq. 2 is approximated with Monte Carlo samples collected from $\pi_{\bm{\theta}}$, and an estimate of $A^{\pi_{\bm{\theta}}}$ is used in place of the true advantages (Schulman et al., 2016). Since off-policy data will bias this gradient estimator, on-policy algorithms discard historic off-policy data after each policy update and collect new data with the updated policy. [^2] In practice, policy gradient algorithms have other sources of gradient bias, such as using $\lambda$-returns in place of Monte Carlo returns (Sutton and Barto, 2018) and using $d^{\gamma}_{\pi_{\bm{\theta}}}$ in place of $d_{\pi_{\bm{\theta}}}$ (Nota and Thomas, 2019).
@@
This foundational idea of policy learning via stochastic gradient ascent was first proposed by Williams (1992) under the name reinforce. Since then, a large body of research has focused on developing more scalable policy gradient methods (Kakade, 2001; Schulman et al., 2015; Mnih et al., 2016; Espeholt et al., 2018; Lillicrap et al., 2015; Haarnoja et al., 2018). Currently, the most successful variant of on-policy learning is proximal policy optimization (PPO) (Schulman et al., 2017), the algorithm of choice in several high-profile success stories (Berner et al., 2019; Akkaya et al., 2019; Vinyals et al., 2019). PPO maximizes the clipped surrogate objective
@@
$\begin{split}{\mathcal{L}}_{\text{PPO}}({\bm{s}},{\bm{a}},{\bm{\theta}},{\bm{\theta}}_{\text{old}})=\min(&g({\bm{s}},{\bm{a}},{\bm{\theta}},{\bm{\theta}}_{\text{old}})A^{\pi_{{\bm{\theta}}_{\text{old}}}}({\bm{s}},{\bm{a}}),\\ &\mathtt{clip}(g({\bm{s}},{\bm{a}},{\bm{\theta}},{\bm{\theta}}_{\text{old}}),1-\epsilon,1+\epsilon)A^{\pi_{{\bm{\theta}}_{\text{old}}}}({\bm{s}},{\bm{a}})),\end{split}\vskip-5.0pt$ (3)
@@
where ${\bm{\theta}}_{\text{old}}$ denotes the policy parameters prior to the update, $g({\bm{s}},{\bm{a}},{\bm{\theta}},{\bm{\theta}}_{\text{old}})$ is the policy ratio $g({\bm{s}},{\bm{a}},{\bm{\theta}},{\bm{\theta}}_{\text{old}})=\frac{\pi_{\bm{\theta}}({\bm{a}}|{\bm{s}})}{\pi_{{\bm{\theta}}_{\text{old}}}({\bm{a}}|{\bm{s}})}$, and the $\mathtt{clip}$ function with hyperparameter $\epsilon$ clips $g({\bm{s}},{\bm{a}},{\bm{\theta}},{\bm{\theta}}_{\text{old}})$ to the interval $1-\epsilon,1+\epsilon]$. This objective disincentivizes large changes to $\pi_{\bm{\theta}}({\bm{a}}|{\bm{s}})$. While other policy gradient algorithms perform a single gradient update per data sample to avoid destructively large policy updates, PPO’s clipping mechanism permits multiple epochs of minibatch policy updates.